I need assistance writing slope-intercept form of a line between two given points. An example would be writing slope-intercept form of a line through points (1,-3) and (0,4.
Answer by Alan3354(31518)

writing slope-intercept form of a line through points (1,-3) and (0,4).
-------------
Find the slope, m
m = diffy/diffx = 7/(-1)
m = -7
------
Use y = mx + b and either point to find b, the y-intercept
-3 = -7*1 + b
b = 4
--> y = -7x + 4
===================
If you know determinants:
|x +y 1|
|1 -3 1| = 0
|0 +4 1|
------------
(Plus signs for alignment)
x*(-3 - 4) - y*(1 - 0) + 1*(4 - 0) = 0
-7x - y + 4 = 0
y = -7x + 4